OpenStudy (anonymous):

Will Fan & Medal! Please help! I think it's D The measures of a pair of complementary angles must sum to 90o . If the angle measures are 3x and 2x -50, what is the value for x? A. 50 B. 90 C. 28 D. 35

OpenStudy (anonymous):

@jhonyy9 Could u help me please?

jhonyy9 (jhonyy9):

so there are angles measured 3x and 2x-50 so what is the value of x ?

jhonyy9 (jhonyy9):

then 3x +2x-50 =90 5x=140 x=140/5 = 28 x=28 ok ?

jhonyy9 (jhonyy9):

C . is right sure

OpenStudy (anonymous):

Okay thank you! Could you help me with 4 more?